OpenStudy (anonymous):

please help me :( will give medal and fan Find the exact circumference. r = 3 1/2 ft A. 14π ft B. 7π ft C. 3.5π ft D. 1.75π ft

OpenStudy (anonymous):

is r=3(1/2) or what... clear this...

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):



OpenStudy (anonymous):

3 1/2=7/3 circumference=2*pi*r =2*pi*7/3

OpenStudy (anonymous):

so is it 2 x 3.14 x 7/3

OpenStudy (anonymous):

3 1/2=7/2 not 7/3 now cir=2*pi*r cir=2*pi*7/2 cir=7pi feet :-) @SilverOwl

OpenStudy (anonymous):

@SilverOwl now my medal mam :-)

OpenStudy (anonymous):

ahh thanks for the medal :-)

OpenStudy (anonymous):

np

OpenStudy (anonymous):

:-)